compute the de broglie wavelength  of electron  in bohrs first orbit ?

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
from standard info . its clear that the velocity of electron in bohr frst orbit = 2.18 * 10 ^6 m / s 
now de broglie wavelength is given by = lambda = h/ mv 
h = planks constant = 6.62 * 10 ^ -34 
m = mass of an electron = 9.1 * 10 ^ -31 
v = velocity of an elecetron 
using this lambda = 3.33 * 10 ^ 10 m = 3.33 armstrong 
HOPE IT CLEARS 
ALL THE BEST ..